# OrionNav: Online Planning for Robot Autonomy with Context-Aware LLM and Open-Vocabulary Semantic Scene Graphs

### v1

@@ IV-C1 Navigation
The ROS2 navigation stack is used to execute the goto commands from the LLM planner. For the global trajectory planning, we use a wavefront Dijkstra search method, that uses a global cost map computed from the occupancy grid map generated using the LiDAR data. The obstacles in the map are inflated radially, with manually tuned inflation radius and cost scaling factor that determine the motion planning penalty around obstacles. Considering the non-linearity in the considered motion model and the importance of maintaining a planning horizon for improved safety, a sampling-based Model Predictive Path Integral (MPPI) controller is used which outputs the desired velocity commands to the locomotion controller on the robot.
@@ IV-C2 CBF-QP Action Filter
Even though the output of the MPPI controller is aware of the local obstacles through the cost map, it does not guarantee the safety of the actions in terms of colliding with surrounding obstacles. To further improve the safety features of our motion planning pipeline, we adopt the formulation proposed in [103] to introduce a safety filter between the output of the MPPI and the input of the low-level velocity tracker on the robot. This low-level filter operates analogous to the reactive instincts of biological animals, creating a low-latency link between the perception (raw LiDAR pointclouds) and action to filter out dangerous commands that lead to collision.
@@ IV-C3 Exploration
To enable efficient exploration, we use the m-explore ROS2 package [^1] https://github.com/robo-friends/m-explore-ros2 based on [104] with some key modifications. The robot’s current 2D position and the occupancy map generated using SLAM (Section IV-A2) are used. The map has grid points classified as unexplored, occupied, or explored. Unexplored grid points adjacent to explored points are clustered, with the centroid of each cluster defined as a frontier point. Each frontier $\mathcal{F}$ is, therefore, described by its 2D coordinates $(\mathcal{F}_{x},\mathcal{F}_{y})$, distance from the robot $\mathcal{F}_{d}$, number of points $\mathcal{F}_{n}$, and a weight $\mathcal{F}_{w}=\alpha\mathcal{F}_{d}+\mathcal{F}_{n}$. We set $\alpha=50$ to prioritize closer frontiers. The highest-weighted frontier is selected for navigation, and non-navigable frontiers, such as the ones behind obstacles or in narrow spaces, are blacklisted. Unlike the original implementation of m-explore, which halts exploration when all frontiers are either explored or blacklisted, we continue exploration by assigning random goals within already explored areas. This strategy allows the robot to detect new frontiers as it moves closer to frontiers that were previously blacklisted due to measurement errors, making the algorithm more robust.
@@ V-A Experiment Setup
The primary goals of our experiments are to demonstrate: (a) the real-world applicability of our pipeline through various end-to-end experiments, (b) the advantages of utilizing a Large Language Model (LLM) as a high-level planner, and (c) the benefits of incorporating hierarchical scene graphs and their integration with the LLM. To achieve these objectives, we structured the experiments into three scenarios: Scenarios A and B focus on object retrieval, where the system locates and navigates to a specific object, and Scenario C addresses room navigation, which involves detecting and navigating to a particular room. We conducted a total of 96 end-to-end experiments, with the detailed setup presented in Table I. In Scenario A (short-range object retrieval), objects were positioned within the same room or area as the robot’s starting point, with no occlusions between the object and the robot. Scenario B (long-range object retrieval) required navigating through corridors and multiple rooms to locate target objects. Additionally, to highlight the benefits of the LLM planner and LLM-based room detection, we performed comparative analyses with baseline methods, as discussed in Section V-C1 and Section V-C2.
Figure 5: Robot Setup: All experiments are conducted using a Unitree Go2 quadrupedal robot equipped with onboard LiDAR, Stereo Camera, and compact computers.
@@
All experiments were conducted using a Unitree Go2 quadruped robot as the navigation agent, as depicted in Fig. 5. This system is equipped with a Hesai LiDAR and a Realsense D455 stereo camera mounted on top of the robot, alongside an NVIDIA Jetson Orin NX and an NVIDIA Jetson Orin Nano for efficient and lightweight computation. Additionally, the quadruped robot utilizes its legged-inertial sensors to aid in the estimation of its odometry.
The experiments were carried out in four office environments featuring multiple rooms and corridors as shown in Fig. 6. These environments contain a variety of typical office objects such as computers, tables, chairs, bags, and potted plants. Furthermore, the environments include various room types such as offices, break rooms, and conference rooms.
Figure 7: Success and failure cases of OrionNav across all experiments. Failure cases are categorized into mapping, exploration, perception, and navigation failures, with detailed breakdowns provided on the right.
Figure 8: Scenario A Examples: The robot observed the queried objects during the first 360-degree rotation and then navigated toward them.
Figure 9: Scenario B Example 1: The system is tasked with locating a bag, which is not visible from the robot’s initial position. The LLM planner issued an exploration command, prompting the robot to explore until the bag is detected, after which the robot successfully navigated to the target.
Figure 10: Scenario B Example 2: The system is tasked with locating a bag. It begins at the starting position (a) and initially searches the office-like space (b). After failing to find the bag, the system initiates global exploration (c, d, e). Upon detecting another office space (f), the system enters and searches it (g), successfully locating the bag.
@@ V-B Results
We present the overall results, including individual failure cases and their corresponding reasons, in Fig. 7. Across a total of 96 trials, we achieved 85 successes, resulting in an overall success rate of over 88%. In two trials, failures occurred due to erroneous SLAM maps, and in one case, a high odometry error caused the system to fail. Excluding these, out of the remaining 93 trials, the exploration algorithm failed to direct the robot to the location containing the queried objects in three cases. From the 90 trials with successful exploration, four trials experienced semantic segmentation errors, where either the object was undetected or misclassified. Additionally, there was one case of failure in the CBF-based navigation system, where the robot collided with an obstacle. Notably, the system did not encounter any failures related to LLM planning or scene graph generation, underscoring their robustness. A detailed breakdown of the success rates for various scenarios is provided in Tab. II, along with the average number of LLM Planner steps. The following subsections offer an in-depth analysis of the system’s performance under different conditions.
@@
Figure 11: Scenario C Example: The system is tasked to go to an office. The LLM planner issued an exploration command, during which, the robot first encountered a break room. The robot continued exploring until it reached the office, at which point the exploration was halted by the LLM planner, successfully completing the task.
V-B1 Scenario A
In this scenario, both the robot and the target are placed in the same region without any occlusions between them, though the robot is often initially oriented away from the target. We conducted 18 experiments in total. Two objects were selected and placed in three different locations within the environment each, resulting in six distinct target locations. For each target location, we performed three experiments starting from similar positions to account for the stochastic nature of the LLM’s outputs. We observed a high success rate of about 92% in this scenario. During the robot’s initial 360-degree rotation, our framework was able to detect most objects in the environment and assign room labels. If the target object was observed during this rotation, the LLM planner issued the goto command, allowing the robot to navigate directly to the object. In cases where the object was not observed during the first pass, the LLM typically issued a search room command, upon which the object was usually detected.
Three examples of this scenario are demonstrated in Fig. 8, where the system is tasked with finding a bag, a potted plant, and a bottle, respectively. In each case, the system successfully located the queried object during the first 360-degree scan and then navigated toward it. Notably, in all instances, the robot was initially oriented away from the target object.
V-B2 Scenario B
In this scenario, the target object was placed at a larger distance from the robot, often located in separate rooms or corridors. Since the robot lacks prior knowledge of the environment, this scenario demands real-time exploration, making it more challenging than Scenario A. We selected different objects and positioned them at various locations, resulting in a total of 8 distinct target locations. As in Scenario A, we performed three experiments starting from similar positions for each target location to account for the stochastic nature of the LLM’s outputs. Despite the increased complexity, we observed a high success rate of 88%. Unlike in Scenario A, the LLM planner issued global exploration commands at the start. By following these exploration commands, the robot successfully navigated the environment, ultimately detecting the target object. Upon detection, the LLM would issue a goto command, guiding the robot to the target location.
In Fig. 9, we present an example of the robot operating in Scenario B, where the system is tasked with finding a bag. Initially, the bag was not visible from the robot’s starting position. After completing a 360-degree scan, the LLM planner issued a global exploration command, as shown in Fig. 9b. The robot then followed the exploration frontiers and eventually detected the bag upon reaching the position illustrated in Fig. 9c. At this point, the LLM planner issued a goto command, and the robot successfully navigated to the bag. Another example of Scenario B is shown in Fig. 10. This is conducted in the larger ‘Environment B’ from Fig. 6.
V-B3 Scenario C
Scenario C focuses on locating specific rooms, such as an office or a break room. As described in Section IV-A3, our system identifies these rooms using the LLM-based room detection module, which relies on the objects present within the rooms. For this set of experiments, we selected two target rooms and initialized the robot from two different positions for each room. Each experiment was repeated three times, with slight variations in the robot’s starting position. In total, we conducted 12 trials in Scenario C, resulting in only 3 failure cases.
A real-world example of Scenario C is illustrated in Fig. 11, where the system is tasked with finding an office room. Initially, the office room is not visible from the robot’s starting position. As a result, the LLM planner issued an exploration command (Fig. 11b). During exploration, the robot encountered a room containing a coffee maker, microwave, refrigerator, table, and several chairs. The room detection module classified this as a break room, as shown in Fig. 11c. Since the target had not been found yet, the LLM planner instructed the system to continue exploring. The robot then entered a room equipped with multiple computers, monitors, tables, cabinets, and chairs, as depicted in Fig. 11d. Upon identifying this room as an office, the system halted exploration, successfully completing the task.
@@ V-B3 Scenario C
V-C Comparative Studies
In this subsection, we compare baseline methods and evaluate the design choices made in the proposed OrionNav, specifically in scenarios involving dynamic environmental changes. The robot is tasked with searching for an object that was initially absent from the environment but added later. For example, this can happen when the object is placed in a room after the robot has already mapped the area. To assess performance, we designed two scenarios: one with a pre-built map (warm start) and another where the robot builds the map from scratch (cold start). In the warm start scenario, the queried object is absent from the pre-built map, as it was placed after mapping was completed. In the cold start scenario, the object was added after the robot had already mapped that specific area. The goal in both cases is to locate the object while minimizing the time required to complete the task. We conducted a total of 9 trials to evaluate our method.
For comparison, we considered two baselines. The first, Frontier-Search, explores the environment by searching through frontiers and transitions to random search once all frontiers are exhausted. This method leverages the ROS2 exploration package, as described in Section IV-C3. To verify object detection in this baseline, the operator manually inspects the object within the object map. The second baseline, Object-Map-Search, utilizes a semantic object map without room labels to locate the target object. For this approach, we introduced a new action primitive, search_object(<object name >), which navigates to the specified object and performs a 360-degree scan to search for other nearby objects. The prompts for the LLM planner were adjusted accordingly.
Figure 13: Example: Dynamically Changes in Environment: The system is provided with a partial map of the environment containing an office room among other locations. A monitor is later placed in the office room and the system is tasked with finding it. In (a), using only the semantic object map without room labels, the system searched near the table in the break room but did not find the monitor. It then explored the environment and eventually located it in the office. In (b), with the proposed hierarchical scene graph, the system identified the office and directly found the monitor inside it.
The comparative results are presented in Fig. 12, where the task completion times are shown as bars for each experiment. Tasks exceeding 10 minutes were deemed failures, indicated by blank bars in the Fig. 12.

### v2

@@ IV-C1 Navigation
The ROS2 navigation stack is used to execute the goto commands generated by the LLM planner. The agent begins by retrieving the locations of the specified objects or rooms from the scene graph. This scene graph is continuously updated as the robot explores and gathers new sensory data. The retrieved location is used to generate a target goal location for global trajectory planning. A cost-aware $A^{*}$ search method is used for global trajectory planning. This method calculates the optimal path to the target location by searching through a grid-based representation of the environment. The global cost map required for this search is computed from an occupancy grid map generated from LiDAR mapping.
To enhance safety and navigation efficiency, obstacles in the occupancy grid map are inflated radially. This inflation accounts for the robot’s physical dimensions and adds a safety buffer to prevent collisions. The inflation radius and cost scaling factors are manually tuned parameters that define how close the robot can approach obstacles and how the proximity to obstacles affects the path cost. Considering the non-linear dynamics inherent in the robot’s motion model and the necessity of maintaining a forward-looking planning horizon for improved safety, we integrate a sampling-based Model Predictive Path Integral (MPPI) controller for local trajectory execution. The MPPI controller outputs the desired velocity commands to the locomotion controller on the robot.
@@ IV-C2 CBF-QP Action Filter
While the MPPI controller accounts for local obstacles through the cost map, it does not inherently guarantee collision-free actions with surrounding obstacles. The MPPI optimizes trajectories based on a predictive model and cost functions but may not react swiftly enough to sudden changes or unmodeled dynamics that could lead to unsafe situations. To further enhance the safety of our motion planning pipeline, we incorporate a Control Barrier Function Quadratic Programming (CBF-QP) safety filter as proposed in [105]. This safety filter is positioned between the MPPI controller’s output and the robot’s low-level velocity tracker, acting as an additional layer of safety oversight.
The CBF-QP safety filter operates analogously to the reactive instincts observed in biological organisms, establishing a low-latency connection between perception and action. Specifically, it directly utilizes raw LiDAR point cloud data to assess the immediate environment and modifies the control commands to avoid collisions proactively. Integrating this filter into our motion planning pipeline significantly enhances the robot’s ability to navigate safely in complex and dynamic environments. It provides a robust safeguard against potential collisions, complementing the predictive capabilities of the MPPI controller with real-time reactive adjustments. This strategy ensures that the robot not only follows optimal paths but also adapts instantaneously to unforeseen obstacles, thereby improving overall navigation safety and reliability.
@@ IV-C3 Exploration
To facilitate efficient and comprehensive exploration of unknown environments, we utilize the m-explore ROS2 package [^1] https://github.com/robo-friends/m-explore-ros2, which is based on the exploration strategy described in [106]. We have implemented several key modifications to adapt the algorithm to our specific requirements and to enhance its robustness in practical applications.
The exploration process relies on the robot’s current 2D position and the occupancy grid map generated using SLAM, as detailed in Section IV-A2. The occupancy grid map represents the environment in a discretized form, where each grid cell is classified into one of three categories based on sensor data: unexplored, occupied, or explored. Unexplored cells are those that the robot has not yet observed, occupied cells contain obstacles or impassable terrain, and explored cells represent free space that has been observed and mapped. To identify areas on the boundary between known and unknown space, we cluster unexplored grid cells that are adjacent to explored cells. This clustering process groups contiguous unexplored cells into distinct regions, and the centroid of each cluster is designated as a frontier point. Frontiers represent gateways to unexplored regions and are critical targets for the robot to expand its knowledge of the environment.
Each frontier $\mathcal{F}$ is characterized by several key parameters, including its 2D coordinates $(\mathcal{F}_{x},\mathcal{F}_{y})$ in the map reference frame, the Euclidean distance $\mathcal{F}_{d}$ from the robot’s current position to the frontier point, and the number of grid cells $\mathcal{F}_{n}$ in the cluster, which indicates the potential area that can be explored by visiting that frontier. We compute a weight $\mathcal{F}_{w}=\alpha\mathcal{F}_{d}+\mathcal{F}_{n}$ for each frontier to prioritize them for exploration. We set $\alpha=50$ to prioritize closer frontiers and balance the trade-off between minimizing travel distance and maximizing exploration gain. A higher value of $\alpha$ increases the influence of proximity, encouraging the robot to select closer frontiers to reduce navigation time and energy consumption, while the term $\mathcal{F}_{n}$ promotes frontiers that lead to larger unexplored areas, maximizing the informational gain.
The robot selects the frontier with the highest weight $\mathcal{F}_{w}$ as the next navigation goal, ensuring efficient expansion of the map by choosing frontiers that offer the best trade-off between proximity and exploration potential. However, not all frontiers are navigable due to environmental constraints such as obstacles blocking the path, narrow passages that the robot cannot traverse, or inaccuracies in the map stemming from sensor noise or localization errors. These non-navigable frontiers are identified through path planning attempts; if the robot cannot find a viable path to a frontier using its navigation stack, the frontier is marked as blacklisted and excluded from future consideration. Unlike the original implementation of m-explore, which terminates the exploration process when all frontiers are either explored or blacklisted, our modified algorithm continues exploration by assigning random goals within the already explored free space when no navigable frontiers are available. This strategy serves multiple important functions. First, by moving to different locations within the explored area, the robot may gain new perspectives on previously blacklisted frontiers. Changes in sensor viewpoints or accumulated mapping data can reveal new paths or correct mapping errors, potentially converting blacklisted frontiers into navigable ones. Second, the robot may uncover new frontiers that were not previously detected due to occlusions, limited sensor range, or initial mapping inaccuracies. Continuous movement increases the likelihood of observing unexplored areas that were not visible from previous positions. Third, revisiting explored areas allows the robot to refine its occupancy grid map, correcting any errors or inconsistencies caused by sensor noise or environmental changes, leading to a more accurate and reliable representation of the environment. By integrating this strategy, we enhance the robustness of the exploration algorithm, ensuring that the robot continues to make progress even when initial exploration efforts stall. This approach mitigates the impact of measurement errors and uncertainties inherent in real-world environments. The robot remains active, persistently seeking opportunities to expand its map and improve its understanding of the surroundings.
In summary, our modifications to the m-explore package enhance autonomous exploration in scenarios involving measurement errors in mapping due to perceptual limitations or error, changes in environment, and complex obstacle configurations. The improved algorithm balances exploration efficiency and thoroughness through weighted frontier selection, frontier blacklisting, and a novel strategy of random goal assignment in explored space when viable frontiers are unavailable. This approach enables continuous exploration even when initial efforts stall due to environmental constraints or mapping inaccuracies. The enhanced capabilities enables the construction of accurate scene graphs to empower the LLM planner to generate more informed and contextually appropriate commands.
@@ V-A Experiment Setup
Our experiments aim to achieve three primary objectives: (a) demonstrate the real-world applicability of our pipeline through various end-to-end experiments (Section V-B), (b) showcase the advantages of utilizing an LLM as a high-level planner (Section V-C1), and (c) highlight the benefits of incorporating hierarchical scene graphs and their integration with the LLM planner (Section V-C2). We organized the experiments into three scenarios to evaluate different aspects of our system. Scenarios A and B focus on object retrieval tasks, where the system must locate and navigate to a specific object. Scenario C addresses room navigation tasks, involving detecting and navigating to a particular room within the environment. In total, we conducted 96 end-to-end experiments, as detailed in Table I. In Scenario A (short-range object retrieval), the target objects were placed within the same room or area as the robot’s starting position, with no occlusions between the robot and the object. This scenario tests the system’s ability to quickly identify and approach objects in an environment with minimal navigation complexity and visual obstructions. Scenario B (long-range object retrieval) presented a more complex challenge, requiring the robot to navigate through corridors and multiple rooms to locate the target objects, thereby testing the system’s mapping, navigation and planning capabilities in a more realistic setting. Scenario C (room navigation) assessed the robot’s ability to interpret higher-level navigation goals by detecting and moving to a specified room, independent of specific objects. This highlights the integration of hierarchical scene understanding with LLM planning.
Figure 5: Robot Setup: OrionNav’s capabilities are demonstrated on a Unitree Go2 quadrupedal robot equipped with onboard LiDAR sensor, stereo camera, and embedded computers equipped with a graphics processing unit (GPU).
@@
The experiments were conducted in four distinct office environments, each featuring multiple rooms and corridors, as depicted in Fig. 5. These environments contain a variety of objects, such as computers, chairs, microwaves, coffee makers, and potted plants, alongside different room types including offices, break rooms, and conference rooms, providing diverse scenarios for the robot’s tasks. All experiments utilized the Unitree Go2 quadruped robot as the physical agent, depicted in Fig. 5. It is equipped with a top-mounted Hesai LiDAR for high-resolution 3D mapping and a Realsense D455 stereo camera for visual perception. All modules in our framework, except for the LLM planner, run onboard the robot, while the LLM planner (GPT-4-Turbo [41]) is accessed via a cloud API. Onboard computation is handled by two key modules: the NVIDIA Jetson AGX Orin (32 GB RAM) and the Jetson Orin Nano (8 GB RAM). The AGX Orin handles computationally intensive tasks like open-vocabulary semantic segmentation, semantic object mapping, and scene graph generation. The Jetson Orin Nano handles mapping and odometry at 10 Hz, matching the LiDAR’s update rate, by utilizing data from LiDAR and legged-inertial sensors and employing the SLAM Toolbox [101]. It also handles the low-level control for navigation, comprising the ROS2 navigation stack, MPPI controller with CBF-QP action filter, and exploration module.
Figure 7: Success and failure cases of OrionNav across all experiments. Visualization of our task execution framework with breakdown of different obsereved failures that occur due to perception, and navigation failures. A detailed breakdown of failures provided on the right.
During the experiments, the stereo camera outputs color and depth frames at 30 Hz, each with a resolution of 640$\times$480. The open-vocabulary semantic segmentation model uses the color frames without resizing to generate object masks at around 10 Hz. The model was optimized for memory and computational efficiency through vectorized operations and TensorRT with BF16 quantization, enabling it to recognize 560 distinct object classes and process 1306 text categories, with the flexibility to dynamically add new classes as needed. The semantic object map is refreshed at 2 Hz by integrating the odometry output, depth frames from the stereo camera, semantic information from the segmentation model, and pose graph corrections from bundle adjustment. The hierarchical scene graph is generated from the semantic object map using spatial clustering and room hierarchies, updating every two seconds. Room label predictions are handled by a quantized LLaMA 3 model, with system prompts containing 21 tokens and user prompts averaging 175 tokens, depending on the input. The LLaMA 3 model processes these predictions at a rate of about 24 tokens per second. For the LLM planner, system prompts comprises 566 tokens, while user prompts averages around 150 tokens.
@@ V-B Results
We visualize the results of all the experiments conducted on the robot using OrionNav in Fig. 7. It presents the overall results of our experiments, including individual failure cases and their corresponding causes. Out of a total of 96 trials, OrionNav successfully completed 85, resulting in a success rate exceeding 88%. Failures due to erroneous SLAM maps occurred in two trials, and a high odometry error caused failure in one trial. Excluding these three instances, the remaining 93 trials proceeded without localization issues. In this subset, the exploration algorithm failed to guide the robot to the location containing the queried objects in three cases. Among the 90 trials with successful exploration, four experienced semantic segmentation errors, where the target object was either undetected or misclassified by the perception system. Additionally, there was one instance of failure in the CBF-based navigation system, resulting in the robot colliding with an obstacle. Importantly, no failures were attributed to the LLM planner or scene graph generator, highlighting their robustness and reliability. A detailed breakdown of the success rates across various scenarios along with the average number of steps taken by the LLM planner is provided in Tab. II. Next, we offer an in-depth analysis of the system’s performance under varying conditions for three different task types, shedding light on the strengths and limitations observed during the experiments.
Figure 8: Short range object navigation task: The robot observes the queried objects during the first 360 rotation and then navigates toward them.
Figure 9: Long range object navigation task: The system is tasked with locating a bag in a corridor environment. Initially, the bag is not visible, prompting the LLM planner to issue an exploration command. The robot explores until it detects the bag, then successfully navigates to it.
Figure 10: Long range object navigation task: The system is tasked with locating a bag on a floor of a building. It begins at the starting position (a) and initially searches the office-like space (b). After failing to find the bag, the system initiates global exploration (c, d, e). Upon detecting another office space (f), the system enters and searches it (g), successfully locating the bag.
V-B1 Scenario A
The robot and the target object are placed within the same region without occlusions between them. We conducted 36 experiments to evaluate our framework’s performance, using four distinct objects placed strategically in various locations. To account for the stochastic nature of the LLM’s outputs, we perform three trials for each target location, starting from the same initial positions. Our framework achieves a high success rate of approximately 88% in tasks related to this scenario. During the robot’s initial 360 scan, the system is able to detect most objects in the environment and assign room labels. If the target object is observed during this initial scan, the LLM planner promptly issues a goto command, allowing the robot to navigate directly to the object. If the target object is not detected during the initial scan, the LLM planner issues a search_room command. This adaptive approach is effective, as the target object is usually detected during the subsequent search phase, demonstrating the robustness of our exploration strategy. Three examples of this scenario are demonstrated in Fig. 8, where the system is tasked with finding a bag, a potted plant, and a bottle, respectively. In each case, the system is successfully able to locate the queried object during the 360 LiDAR scan and then navigate towards, away from the target object.
V-B2 Scenario B
[…]
Figure 11: Room navigation task: The system is tasked with locating an office room on a building floor. The LLM planner initiates the task by issuing an exploration command, during which the robot first encounters a break room. Continuing its search, the robot eventually reaches an office room which is correctly detected and classified during scene graph generation. The LLM planner then halts exploration, successfully completing the task.
V-B3 Scenario C
Scenario C focuses on locating specific rooms, such as offices or break rooms. As described in Section IV-A3, our system identifies rooms using the LLM-based room detection module, which relies on the objects present within each room. For this set of experiments, we select four target rooms and initialize the robot from different positions. Each experiment is repeated three times with slight variations in the robot’s starting position, resulting in a total of 15 trials and 12 successes. Fig. 11 illustrates a real-world example of Scenario C, where the system is tasked with finding an office room that is not initially visible from the robot’s starting position. Consequently, the LLM planner issues an exploration command (Fig. 11b). During exploration, the robot encounters a room containing a coffee maker, microwave, refrigerator, table, and several chairs. The room detection module classifies this as a break room (Fig. 11c). Since the target has not been found, the LLM planner instructs the robot to continue exploring. The robot then enters a room equipped with multiple computers, monitors, tables, cabinets, and chairs (Fig. 11d). Upon identifying this room as an office, the system halts exploration, successfully completing the task.
@@ V-B3 Scenario C
V-C Comparative Studies
We compare baseline methods and evaluate the design choices of the proposed OrionNav system, particularly in scenarios involving dynamic environmental changes. In each experiment, the robot is tasked with searching for an object that is initially absent from the environment and introduced at a later stage, for e.g., when an object is placed in a room after the robot has already mapped the area. To assess performance, we design two scenarios: one using a pre-built map and another where the robot builds the map from scratch. In both scenarios, the queried object is added to the environment after the mapping phase is complete. These setups allow us to evaluate the system’s ability to adapt to changes in the environment, both with and without prior knowledge of the map. The goal is to locate the object while minimizing the time required to complete the task. We conduct nine trials to evaluate our method and compare it with baseline approaches.
We consider Frontier-Search and Object-Map-Search baselines to demonstrate the efficacy of our framework. Frontier-Search explores the environment by searching through frontiers and transitions to random search once all frontiers are exhausted. This method leverages the ROS2 exploration package, as described in Section IV-C3. Since this baseline lacks an integrated mechanism to recognize when the target object has been detected, an operator manually inspects the generated object map for the presence of the queried object after the exploration concludes. The second baseline, Object-Map-Search, utilizes a semantic object map without room labels to locate the target object. A new action primitive, search_object(<object name >) is introduced for this task. It instructs the robot to navigate to the specified object in the map and perform a 360 scan to search for other nearby objects. Fig. 12 presents the comparative results of task completion times for each experiment as bars, with tasks classified as failures when exceeding 10 minutes.
@@ V-C Comparative Studies
Figure 13: Autonomous object navigation in changing environment: The system is provided with a partial map of the environment containing an office room among other locations. A monitor is later placed in the office room and the system is tasked with finding it. In (a), using only the semantic object map without room labels, the system searched near the table in the break room but did not find the monitor. It then explored the environment and eventually located it in the office. In (b), with the proposed hierarchical scene graph, the system identified the office and directly found the monitor inside it.